OpenStudy (anonymous):

solve: 3(x+2)-2=-(5+x)+x how do I solve this problem on a scientific calculator?

OpenStudy (campbell_st):

well you can't you would have to do it on paper... distribute both sides 2x + 6 - 2 = -5 -x + x simplify 2x + 4 = -5 now just solve for x

hero (hero):

Depends on what kind of calculator you have actually.

OpenStudy (precal):

you can solve it using a graphing calculator but not a scientific calculator